OpenStudy (anonymous):

Solve, using the square root method: 5x2 = 60. Round the solution to the nearest tenth. A. {3.5, -3.5} B. {7.7, -7.7} C. {3.5} D. No real solution

OpenStudy (anonymous):

start by dividing both sides by 5

OpenStudy (anonymous):

then tell me what did you get...

OpenStudy (anonymous):

\[5\div60 \]

OpenStudy (anonymous):

=12

OpenStudy (anonymous):

X^2=12 `

OpenStudy (anonymous):

ok... now we have: \[\LARGE x^2=12\] geting square root to both sides we get: \[\LARGE \sqrt{x^2}=\sqrt{12}\] so \[\LARGE x=\pm\sqrt{12}\]

OpenStudy (anonymous):

3.46

OpenStudy (anonymous):

that's right..

OpenStudy (anonymous):

so 3.5

OpenStudy (anonymous):

yes +3.5 and -3.5 \[\LARGE \smile \]

OpenStudy (anonymous):

Ok thank you bye`

OpenStudy (anonymous):

yw bye..